# From Planetesimals to Dwarf Planets by Pebble Accretion

### v1

@@
(February 6, 2025)
The size distribution of trans-Neptunian objects (TNOs) in the Kuiper Belt provides crucial insights into the formation and evolution of the outer Solar System. Recent observational surveys, including OSSOS++, have revealed that dynamically cold and hot TNO populations exhibit similar size distributions for dimmer objects ($H_{r}>5$), which are consistent with planetesimal formation by streaming instability (SI). However, the hot population contains a significantly larger number of massive bodies, including several dwarf planets. In this study, we investigate the role of pebble accretion in shaping the size distribution of hot TNOs, after their formation in the primordial disk (PB) between 20 and 30 au and before these bodies were dynamically implanted into their current orbits by a migrating Neptune. We find that pebble accretion grows the most massive bodies only, consistent with the flattening of the distribution brightwards of $H_{r}=5$. All results point to a correlation (degeneracy) between the pebble aerodynamic size and the intensity of the gas motions. Nevertheless, accretion from an inward-drifting stream of pebbles is unlikely, as it puts extreme demands on the mass budget of pebbles. In particular, the masses of the cold classicals are too low to trigger pebble accretion. Accretion in an environment where pebbles are entrained, as believed to be the case in ALMA rings, is preferable. Combining the constraints obtained from this study with ALMA imagery morphology fitting reveals a typical pebble aerodynamic size of $\tau_{s}\sim 10^{-2}$, a turbulent diffusivity parameter $\alpha_{D}\sim 10^{-3}$, and a total accreted pebble mass of ${\sim}10\,m_{\oplus}$ in the primordial belt. Those TNOs formed through significant pebble accretion with masses exceeding ${\sim}10^{-4}\,m_{\oplus}$ are likely to satisfy the International Astronomical Union’s “round shape” criterion for dwarf planets.
@@ 1 Introduction
The trans-Neptunian space, conventionally known as the Kuiper belt, hosts numerous icy bodies which are believed to be ancient remnants of the solar system’s outer planetesimal disk. As more trans-Neptunian objects (TNOs) are discovered by various modern observational surveys, such as the Canada–France Ecliptic Plane Survey (CFEPS; Jones et al. 2006), the Deep Ecliptic Survey (DES; Adams et al. 2014), the Outer Solar System Origins Survey (OSSOS; Bannister et al. 2018), the Dark Energy Survey (DES; Bernardinelli et al. 2022), and the DECam Ecliptic Exploration Project (DEEP; Trilling et al. 2024), there is a growing consensus that the Kuiper belt consists of two distinct populations: dynamically cold TNOs, formed in-situ ($\sim$45 au), and dynamically hot TNOs, which are believed to have formed closer to the Sun ($\sim$20 au), before being implanted to their current locations.
@@
Cold TNOs, also known as cold classicals, are characterized by their low inclinations and relatively stable orbits, suggesting they formed in their current locations and have remained dynamically cold (Petit et al., 2011). In contrast, hot TNOs, which include the hot classicals, resonant objects, and scattering disk objects (see Gladman et al. 2008 for definitions of TNOs dynamical classes), exhibit higher inclinations and eccentricities, suggesting a more violent dynamical history. Moreover, cold TNOs tend to have redder colors, indicating a different surface composition and evolution history compared to the more neutral colors of hot TNOs (Doressoundiram et al., 2002; Trujillo & Brown, 2002; Pike & Lawler, 2017; Schwamb et al., 2019; Müller et al., 2020; Fernández-Valenzuela et al., 2021). The binary fraction is also higher among cold TNOs (Noll et al., 2020), especially those with comparable sizes. Additionally, dwarf planet-sized TNOs are predominantly found in the hot populations, with cold TNOs lacking large objects (Brown, 2008; Fraser et al., 2014). All these differences—orbital characteristics, colors, binary fractions, and size distributions—support the hypothesis that cold and hot TNOs formed at different radial locations and experienced different dynamical evolutions. Cold TNOs likely formed in-situ at their current locations, while hot TNOs formed closer to the Sun and were subsequently implanted into their current orbits during the migration of the giant planets (Malhotra, 1993; Gomes, 2003; Levison et al., 2008; Nesvorný, 2015).
@@
The implantation process of hot TNOs is closely linked to the planetesimal-driven migration of the giant planets, particularly Neptune. In this hypothesis, Neptune is believed to have formed closer to the Sun and migrated outward into a primordial planetesimal belt that spanned $\sim$20–30 au (Gomes, 2003; Levison et al., 2008), with a total estimated mass of ${\sim}20\,m_{\oplus}$ (Nesvorný, 2015). Most of the planetesimals were scattered inward consecutively by Neptune, Uranus, and Saturn, until their orbits cross the orbit of Jupiter. The strong gravity of Jupiter effectively ejected the vast majority of these planetesimals out of the Solar System, leaving only a small fraction to form the Jupiter Trojans (Morbidelli et al., 2005; Nesvorný et al., 2013). Another fraction of the planetesimals were scattered outward by a migrating Neptune into the trans-Neptunian region, forming the hot or implanted population (Malhotra, 1993; Nesvorný et al., 2016).
@@
As the number of discovered TNOs continues to grow, the size distribution [^1] The absolute magnitude measured with the $r$ filter ($H_{r}$) is often used as a handle to study the size of TNOs. For the conversion between $H_{r}$ and TNO mass, see Section 2.1 and Appendix A. of both cold and hot TNOs is becoming increasingly well-constrained. Kavelaars et al. (2021) reported the $H_{r}$ distribution of the cold classical belt from $H_{r}\simeq 5$ to 12 detected by OSSOS, and found an exponentially tapered power-law with an index of $\alpha\approx 0.4$. This functional form is also consistent with the DEEP survey (Napier et al., 2024). Recently, streaming instability (SI) has emerged as one of the leading theories for planetesimal formation in the solar nebula, as evidenced by the high binary fractions observed in cold classicals (Nesvorný et al., 2010, 2019), and the discoveries of contact binaries such as Arrokoth (McKinnon et al., 2020; Lyra et al., 2021). The exponentially tapered power-law is also consistent with SI simulations that model the planetesimal formation process (Schäfer et al., 2017; Abod et al., 2019).
@@
In a follow-up study of the OSSOS++ survey, Petit et al. (2023) conducted a similar analysis for the hot belt. They found that the shape of the $H_{r}$ distribution of the hot belt is similar to that of the cold belt, ranging from $H_{r}\approx 5$ to $H_{r}\approx 8.3$ (corresponding roughly to bodies with $D\approx 90\,\mathrm{km}$ in diameter [^2] The relation between the $H_{r}$ magnitude, mass, and diameter is detailed in Appendix A. When quoting diameters, a nominal internal density of $1\,\mathrm{g\,cm^{-3}}$ is assumed.). This upper limit corresponds to the sensitivity threshold of the OSSOS++ survey at $\sim$48 au, given an apparent magnitude limit of $m_{r}\approx 25$. They also found that the all dynamically hot TNOs and those limited to the main belt exhibit remarkably similar $H_{r}$ distributions (see Figure 3 in Petit et al. 2023), further supporting the shared origin of the dynamically hot TNOs. Given the similarity in the size distributions of the cold and hot TNOs in the low mass end ($H_{r}\gtrsim 5$, roughly $D\gtrsim 400$ km), it is natural to postulate that both populations started off with a similar initial mass function (IMF) given by the SI process. In contrast, the (future) hot population had the high mass end altered by other planet formation processes.
@@
Figure 2: Cumulative number distribution ($N_{<m}$) of bodies in the cold belt and the dynamical hot P48. The cold belt is well described by the analytical fit of Kavelaars et al. (2021) (Equation (2)), whose shape is assumed to be identical to the distribution of the bodies in the primordial belt. Population P48 is reconstructed by taking all bodies brighter than $H_{r}=5$ and within 48 au from the MPC database (histograms) while bodies $5\leq H_{r}\leq 8.3$ are assumed to follow the completeness-corrected distribution from OSSOS++. The corresponding cumulative mass function is given by the black curve. The conversion between mass and magnitude (top) assumes an albedo typical to hot belt bodies (see Appendix A). The dotted lines labeled $p=-2.0$ and $p=-2.5$ correspond to $N_{<m}\propto m^{1+p}$ with $p=-2$ indicating an equal amount of mass per logarithmically spaced mass bin (the corresponding indices in terms of a radius spectrum $dN/dR\propto R^{q}$ are $q=-4$ and $q=-5.5$, respectively). Pebble accretion operates to selectively grow the largest TNOs, flattening the size distribution.
@@
The process that is envisioned is schematically illustrated in Figure 1. SI forms bodies that follow a tapered-exponential in both the cold and the primordial belts. But only the primordial belt featured conditions for pebble accretion to materialize. While the size distribution of smaller bodies stays intact and continuous to follow the initial mass function (IMF) bodies obtain after SI, the larger bodies consume pebbles. Afterwards, the dynamical instability implanted these bodies into the trans-Neptunian space.
Pebble accretion is sensitive to the physical properties of the particles, as well as the local disk conditions. Two parameters in particular are responsible for the characteristics of the accretion: the particle aerodynamic size $\tau_{s}$ and the velocity $\Delta v$ at which pebbles approach the planetary body. In this work, we conduct a quantitative study to verify which combination of these and other parameters are consistent with the size distribution of current TNOs. To this effect, we carefully select a sample of TNOs that are complete, based on the OSSOS++ survey and the Minor Planet Center (MPC) database. For simplicity, we limit the number of parameters in our model (e.g., we only treat one particle size). As one of our key results, we find a tight relation between pebble aerodynamic size $\tau_{s}$ and relative velocity $\Delta v$, which favors formation in an environments that resemble the azimuthally symmetric structures seen in ALMA imagery (Bae et al., 2023), commonly referred to as ”ALMA rings”.
The structure of the Paper is as follows. In Section 2, we describe our simplified model to follow the growth of planetary bodies by pebbles. The simplified nature allows exploration by MCMC methods (Section 3). In Section 4, the implications of these findings are reported in the context of solar system formation models. Further assessment of our model is given in Section 5 and the conclusions are given in Section 6.
@@ 2 Model
This section consists of three parts. In Section 2.1, we reconstruct a population of TNOs that is complete, against which the model outcome can be evaluated. Section 2.2 describes the pebble accretion model, which is formulated in dimensionless units. While this approach makes the model more abstract, it avoids the need of introducing a specific disk model. Finally, in Section 2.3, we construct the MCMC likelihood function and outline the model parameters.
@@ 2.1 The mass function of the dynamical hot population
In Figure 2, we have reconstructed the present-day mass distribution of dynamically hot TNOs within 48 au, a population we refer to as P48. This distribution is composed of two parts. At high masses, we take the bodies directly from the Minor Planet Center (MPC) database [^3] The file distant_extended.dat was retrieved from https://www.minorplanetcenter.net/data on November 9 2024. and filter out the cold classicals by applying an $i_{\text{free}}<4^{\circ}$ cut (Huang et al., 2022). Specifically, we assume the data, which includes the entire hot belt, inner belt, 3:2 population (plutinos), 2:1 population (twotinos), as well as scattering TNOs, is complete for $H_{r}<5$ bodies within $a\leq 48\,\mathrm{au}$ (Weryk et al., 2016; Petit et al., 2023).
We exclude 13 members that belong to the Haumea collisional family, as they are likely fragments ejected from Haumea during/after Neptune’s migration (Proudfoot & Ragozzine, 2019), and are therefore unrelated to the prior pebble accretion process. To account for the lost mass of the Haumea impactor (which could be a moon of Haumea or a dwarf planet from the scattering disk depending on the hypothesis), we include a new body with a mass equal to 3% of Haumea’s (Vilenius et al., 2018; Pike et al., 2020) in our sample. Additionally, we add Neptune’s largest moon, Triton, [^4] This assumes that the implantation probability of bodies ending up as TNOs is similar to the capture probability by Neptune. which is believed to be a captured moon from the primordial belt (Agnor & Hamilton, 2006), as well as Pluto’s largest moon, Charon, which likely formed from a giant impact (Canup, 2005). In total, 66 bodies constitute the high-mass portion of the P48 population (see Table 4), reflecting our current best understanding of the dwarf planet size distribution.
TNOs dimmer than $H_{r}=5$ in P48 are modeled on the debiased distribution of the hot belt [^5] Haumea family members with $H_{r}>5$ are not corrected, as the OSSOS survey reveals a significant scarcity of small members within the Haumea family (Pike et al., 2020)., based on the OSSOS++ survey (Petit et al., 2023). Their Figure 1 presents the differential distribution, which we fit by a broken power law
@@ 2.2 Growth of bodies by pebble accretion
where $\chi_{\bullet}$ is the ratio of the physical-to-Hill radius
@@
where $\rho_{\bullet}$ the internal density of the planetary body and $A_{\mathrm{geo}}=\pi/3^{2/3}\approx 1.5$. Also, the characteristic growth timescale is
@@
where $\rho_{\mathrm{peb}}=Z_{\mathrm{peb}}\rho_{\mathrm{gas}}=Z_{\mathrm{peb}}\Sigma_{\mathrm{gas}}/\sqrt{2\pi}H_{\mathrm{gas}}$ the mass density of pebbles in the midplane, $Z_{\mathrm{peb}}$ the corresponding local “metallicity”, $\Sigma_{\mathrm{gas}}$ the surface density of the gas and $H_{\mathrm{gas}}$ the gas pressure scaleheight. The disk aspect ratio $h=H_{\mathrm{gas}}/r$ can also be written
@@
The degree of coupling between particles and gas is quantified by the particles’ stopping time $t_{\mathrm{stop}}$, which represents the time for a particles to adjust its motion to that of gas. Its dimensionless equivalent, $\tau_{s}=\Omega_{K}t_{\mathrm{stop}}$, is referred to as the aerodynamic size (in the literature, ”Stokes number” or St is also common). In this work, we define pebbles as particles of $\tau_{s}<1$. For Epstein drag, $\tau_{s}$ can be expressed as $\tau_{s}=\pi(\rho_{\bullet}R)_{\mathrm{peb}}/2\Sigma_{\mathrm{gas}}$, indicating that aerodynamic size is proportional to $R_{\mathrm{peb}}$.
The relative velocity $\Delta v=\sigma v_{K}$ can be identified with either:
  • The eccentricity of planetesimals, $\Delta v\simeq ev_{K}$ ($\sigma=e$). This is not considered in this work.
@@
    (e.g., Weidenschilling, 1977) is a parameterization of the radial pressure gradient. Pebbles with $\tau_{s}<1$ follow the gas, while planetesimals ($\tau_{s}\gg 1$) move at the Keplerian velocity. Hence, the pebble-planetesimal velocity is also $\eta v_{K}$.
@@
Crudely, the largest of these velocities – $e$, $\eta$, or $\mathcal{M}h$ – determines $\Delta v$.
Gravitational focusing increases the accretion rate by a factor $(v_{\mathrm{esc}}/\Delta v)^{2}=2Gm/R(\Delta v)^{2}$, where $v_{\mathrm{esc}}$ is the escape velocity from its surface. In dimensionless units, the rate becomes
@@
Gravitational focusing significantly enhances the growth rates of planetesimals. On the other hand, its efficacy is hampered by the more copious non-collisional scatterings of planetesimals – a process known as viscous stirring (Ida & Makino, 1993) – or turbulent stirring (Ida et al., 2008), which increases $\Delta v$. For small pebbles tightly coupled to the gas, this is not a concern, because gas drag circularizes their orbits before the next encounter. Therefore, in contrast to planetesimals, there is no need to follow the evolution of $\Delta v$ with time.
More potently, pebbles are accreting by the settling mechanism – better known as pebble accretion (Ormel & Klahr, 2010; Lambrechts & Johansen, 2012). In pebble accretion, pebbles can be captured at distances much larger than $R$, whereafter the pebbles settle down in the gravitational potential. We adopt the numerically-calibrated expression by Ormel & Liu (2018) [^6] Ormel & Liu (2018) express these expression in terms of an efficiency $\epsilon=\dot{M}/2\pi r\Sigma v_{r}$. We take their 3D expression, $\epsilon=0.39q/\eta h_{\mathrm{peb}}\times f_{\mathrm{set}}^{2}$. With the drift velocity $v_{r}=2\eta v_{K}\tau_{s}$ and $\rho_{\mathrm{peb}}=\Sigma_{\mathrm{peb}}/\sqrt{2\pi}h_{\mathrm{peb}}$, we obtain Equation (10).:
$\frac{dq}{dt}=A_{\mathrm{peb}}\frac{q\tau_{s}}{t_{0}}f_{\mathrm{set}}^{2},$ (10)
with $A_{\mathrm{peb}}=12.3$ and $f_{\mathrm{set}}\leq 1$ being a modulation factor that suppresses accretion rates at small masses. It is a defining characteristic of pebble accretion that rates do not depend on the physical radius (or the nondimensional $\chi_{\bullet}$ parameter; Ormel 2017). In addition, we have assumed the 3D accretion regime for pebbles in Equation (10), for which $\sigma$ does not enter explicitly. The velocity matters in the modulation factor $f_{\mathrm{set}}$, however:
$f_{\mathrm{set}}^{2}=\exp\left[-\left(\frac{\Delta v}{v_{\ast}}\right)^{2}\right]\quad\textrm{or}\quad f_{\mathrm{set}}^{2}=\left[1+a_{\mathrm{turb}}\left(\frac{\Delta v}{v_{\ast}}\right)^{2}\right]^{-3},$ (11)
with $v_{\ast}=(q_{p}/\tau_{s})^{1/3}v_{K}$ and $a_{\mathrm{turb}}=0.33$ (Liu & Ormel, 2018). The first expression in Equation (11) holds for a laminar-dominated velocity field and the second for a turbulent-dominated $\Delta v$. In the case of a turbulent flow, $\Delta v$ should be regarded as the rms value of an underlying velocity distribution that is isotropic (Ormel & Liu, 2018).

### v2

@@
(February 27, 2025)
The size distribution of trans-Neptunian objects (TNOs) in the Kuiper Belt provides crucial insights into the formation and evolution of the outer Solar System. Recent observational surveys, including the Outer Solar System Origins Survey (OSSOS++), have revealed that dynamically cold and hot TNO populations exhibit similar size distributions for dimmer objects ($H_{r}>5$), which are consistent with planetesimal formation by streaming instability (SI). However, the hot population contains a significantly larger number of massive bodies, including several dwarf planets. In this study, we investigate the role of pebble accretion in shaping the size distribution of hot TNOs, after their formation in the primordial disk (PB) between 20 and 30 au and before these bodies were dynamically implanted into their current orbits by a migrating Neptune. We find that pebble accretion grows the most massive bodies only, consistent with the flattening of the distribution brightwards of $H_{r}=5$. All results point to a correlation (degeneracy) between the pebble aerodynamic size and the intensity of the gas motions. Nevertheless, accretion from an inward-drifting stream of pebbles is unlikely, as it puts extreme demands on the mass budget of pebbles. In particular, the masses of the cold classicals are too low to trigger pebble accretion. Accretion in an environment where pebbles are entrained, as believed to be the case in rings seen with the Atacama Large Millimeter Array (ALMA), is preferable. Combining the constraints obtained from this study with ALMA imagery morphology fitting reveals a typical pebble aerodynamic size of $\tau_{s}\sim 10^{-2}$, a turbulent diffusivity parameter $\alpha_{D}\sim 10^{-3}$, and a total accreted pebble mass of ${\sim}10\,m_{\oplus}$ in the primordial belt. Those TNOs formed through significant pebble accretion with masses exceeding ${\sim}10^{-4}\,m_{\oplus}$ are likely to satisfy the International Astronomical Union’s “round shape” criterion for dwarf planets.
@@ 1 Introduction
The trans-Neptunian space, conventionally known as the Kuiper belt, hosts numerous icy bodies which are believed to be ancient remnants of the Solar System’s outer planetesimal disk. As more trans-Neptunian objects (TNOs) are discovered by various modern observational surveys, such as the Canada–France Ecliptic Plane Survey (CFEPS; Jones et al. 2006), the Deep Ecliptic Survey (DES; Adams et al. 2014), the Outer Solar System Origins Survey (OSSOS; Bannister et al. 2018), the Dark Energy Survey (DES; Bernardinelli et al. 2022), and the DECam Ecliptic Exploration Project (DEEP; Trilling et al. 2024), there is a growing consensus that the Kuiper belt consists of two distinct populations: dynamically cold and dynamically hot TNOs.
@@
Cold TNOs, also known as cold classicals, are characterized by their low inclinations and relatively stable orbits, suggesting that they formed in their current locations and have remained dynamically cold (Petit et al., 2011). In contrast, hot TNOs, which include the hot classicals, resonant objects, and scattering disk objects (see Gladman et al. 2008 for the definitions of the TNO dynamical classes), exhibit higher inclinations and eccentricities, suggesting a more violent dynamical history. Moreover, cold TNOs tend to have redder colors, indicating a different surface composition and evolution history compared to the more neutral colors of hot TNOs (Doressoundiram et al., 2002; Trujillo & Brown, 2002; Pike & Lawler, 2017; Schwamb et al., 2019; Müller et al., 2020; Fernández-Valenzuela et al., 2021). The binary fraction is also higher among cold TNOs (Noll et al., 2020), especially those of comparable sizes. Additionally, dwarf planet-sized TNOs are predominantly found in the hot populations, with cold TNOs lacking large objects (Brown, 2008; Fraser et al., 2014). All these differences—orbital characteristics, colors, binary fractions, and size distributions—support the hypothesis that cold and hot TNOs formed at different radial locations and experienced different dynamical evolutions. Cold TNOs likely formed in situ at their current locations ($\sim$45 au), while hot TNOs formed closer to the Sun ($\sim$20 au) and were subsequently implanted into their current orbits during the migration of the giant planets (Malhotra, 1993; Gomes, 2003; Levison et al., 2008; Nesvorný, 2015).
@@
The implantation process of hot TNOs is closely linked to the planetesimal-driven migration of the giant planets, particularly Neptune. In this hypothesis, Neptune is believed to have formed closer to the Sun and migrated outward into a primordial planetesimal belt that spanned $\sim$20–30 au (Gomes, 2003; Levison et al., 2008), with a total estimated mass of ${\sim}20\,m_{\oplus}$ (Nesvorný, 2015). Most of the planetesimals were scattered inward consecutively by Neptune, Uranus, and Saturn, until their orbits crossed the orbit of Jupiter. The strong gravity of Jupiter effectively ejected the vast majority of these planetesimals out of the Solar System, leaving only a small fraction to form the Jupiter trojans (Morbidelli et al., 2005; Nesvorný et al., 2013). Another fraction of the planetesimals were scattered outward by a migrating Neptune into the trans-Neptunian region, forming the hot or implanted population (Malhotra, 1993; Nesvorný et al., 2016).
@@
As the number of discovered TNOs continues to grow, the size distribution [^1] The absolute magnitude measured with the $r$ filter ($H_{r}$) is often used to study the size of TNOs. For the conversion between $H_{r}$ and TNO mass, see Section 2.1 and Appendix A. of both cold and hot TNOs is becoming increasingly well-constrained. Kavelaars et al. (2021) reported the $H_{r}$ distribution of the cold classical belt from $H_{r}\simeq 5$ to 12 detected by OSSOS, and found an exponentially tapered power law with an index of $\alpha\approx 0.4$. This functional form is also consistent with the DEEP survey (Napier et al., 2024). Recently, streaming instability (SI) has emerged as one of the leading theories for planetesimal formation in the solar nebula, as evidenced by the high binary fractions observed in cold classicals (Nesvorný et al., 2010, 2019), and the discoveries of contact binaries such as Arrokoth (McKinnon et al., 2020; Lyra et al., 2021). The exponentially tapered power law is also consistent with SI simulations that model the planetesimal formation process (Schäfer et al., 2017; Abod et al., 2019).
@@
In a follow-up study of the OSSOS++ survey, Petit et al. (2023) conducted a similar analysis for the hot belt. They found that the shape of the $H_{r}$ distribution of the hot belt is similar to that of the cold belt, ranging from $H_{r}\approx 5$ to $H_{r}\approx 8.3$ (corresponding roughly to bodies with $D\approx 90\,\mathrm{km}$ in diameter [^2] The relation between the $H_{r}$ magnitude, mass, and diameter is detailed in Appendix A. When quoting diameters, a nominal internal density of $1\,\mathrm{g\,cm^{-3}}$ is assumed.). This upper limit corresponds to the sensitivity threshold of the OSSOS++ survey at $\sim$48 au, given an apparent magnitude limit of $m_{r}\approx 25$. They also found that all the dynamically hot TNOs and those limited to the main belt exhibit remarkably similar $H_{r}$ distributions (see Figure 3 in Petit et al. 2023), further supporting the shared origin of the dynamically hot TNOs. Given the similarity in the size distributions of the cold and hot TNOs at the low-mass end ($H_{r}\gtrsim 5$, roughly $D\gtrsim 400$ km), it is natural to postulate that the two populations started off with similar initial mass functions (IMF) given by the SI process. In contrast, the (future) hot population had the high-mass end altered by other planet formation processes.
@@
Figure 2: Cumulative number distribution ($N_{<m}$) of bodies in the cold belt and the dynamical hot P48. The cold belt is well described by the analytical fit of Kavelaars et al. (2021) (Equation (2)), whose shape is assumed to be identical to the distribution of the bodies in the primordial belt. Population P48 is reconstructed by taking all bodies brighter than $H_{r}=5$ and within 48 au from the MPC database (histograms) while bodies $5\leq H_{r}\leq 8.3$ are assumed to follow the completeness-corrected distribution from OSSOS++. The corresponding cumulative mass function is given by the black curve. The conversion between mass and magnitude (top) assumes an albedo typical for hot belt bodies (see Appendix A). The dotted lines labeled $p=-2.0$ and $p=-2.5$ correspond to $N_{<m}\propto m^{1+p}$ with $p=-2$ indicating an equal amount of mass per logarithmically spaced mass bin (the corresponding indices in terms of a radius spectrum $dN/dR\propto R^{q}$ are $q=-4$ and $q=-5.5$, respectively). Pebble accretion operates to selectively grow the largest TNOs, flattening the size distribution.
@@
The process that is envisioned is schematically illustrated in Figure 1. SI forms bodies that follow a tapered-exponential in both the cold and the primordial belts, but only the primordial belt featured conditions for pebble accretion to materialize. While the size distribution of smaller bodies stayed intact and continued to follow the initial mass function (IMF) bodies obtained after SI, the larger bodies consumed pebbles. Afterwards, the dynamical instability implanted these bodies into the trans-Neptunian space.
Pebble accretion is sensitive to the physical properties of the particles, and to the local disk conditions. Two parameters in particular are responsible for the characteristics of the accretion: the particle aerodynamic size $\tau_{s}$ and the velocity $\Delta v$ at which pebbles approach the planetary body. For this work, we conducted a quantitative study to verify which combination of these and other parameters are consistent with the size distribution of current TNOs. To this end, we carefully selected a sample of TNOs that are complete, based on the OSSOS++ survey and the Minor Planet Center (MPC) database. For simplicity, we limited the number of parameters in our model (e.g., we only treated one particle size). As one of our key results, we find a tight relation between pebble aerodynamic size $\tau_{s}$ and relative velocity $\Delta v$, which favors formation in an environments that resemble the azimuthally symmetric structures seen in images taken by the Atacama Large Millimeter Array (ALMA; Bae et al. 2023), commonly referred to as ALMA rings.
The structure of the paper is as follows. In Section 2 we describe our simplified model to follow the growth of planetary bodies by pebbles. The simplified nature allows exploration by Markov chain Monte Carlo (MCMC) methods (Section 3). In Section 4, the implications of these findings are reported in the context of formation models for our Solar System. Further assessment of our model is given in Section 5 and the conclusions are given in Section 6.
@@ 2 Model
This section consists of three parts. In Section 2.1, we reconstruct a population of TNOs that is complete, against which the model outcome can be evaluated. Section 2.2 describes the pebble accretion model, which is formulated in dimensionless units. While this approach makes the model more abstract, it avoids the need to introduce a specific disk model. Finally, in Section 2.3, we construct the MCMC likelihood function and outline the model parameters.
@@ 2.1 The mass function of the dynamical hot population
In Figure 2, we have reconstructed the present-day mass distribution of dynamically hot TNOs within 48 au, a population we refer to as P48. This distribution is composed of two parts. At high masses, we take the bodies directly from the Minor Planet Center (MPC) database [^3] The file distant_extended.dat was retrieved from https://www.minorplanetcenter.net/data on November 9 2024. and filter out the cold classicals by applying an $i_{\text{free}}<4^{\circ}$ cut (Huang et al., 2022). Specifically, we assume that the data, which includes the entire hot belt, inner belt, 3:2 population (plutinos), 2:1 population (twotinos), as well as scattering TNOs, is complete for $H_{r}<5$ bodies within $a\leq 48\,\mathrm{au}$ (Weryk et al., 2016; Petit et al., 2023).
We excluded 13 members that belong to the Haumea collisional family, as they are likely fragments ejected from Haumea during or after Neptune’s migration (Proudfoot & Ragozzine, 2019), and are therefore unrelated to the prior pebble accretion process. To account for the lost mass of the Haumea impactor (which could be a moon of Haumea or a dwarf planet from the scattering disk depending on the hypothesis), we include a new body with a mass equal to 3% of Haumea’s (Vilenius et al., 2018; Pike et al., 2020) in our sample. Additionally, we add Neptune’s largest moon, Triton, [^4] This assumes that the implantation probability of bodies ending up as TNOs is similar to the capture probability by Neptune. which is believed to be a captured moon from the primordial belt (Agnor & Hamilton, 2006), as well as Pluto’s largest moon, Charon, which likely formed from a giant impact (Canup, 2005). In total, 66 bodies constitute the high-mass portion of the P48 population (see Table 4), reflecting our current best understanding of the dwarf planet size distribution.
Trans-Neptunian Objects dimmer than $H_{r}=5$ in P48 are modeled on the debiased distribution of the hot belt [^5] Haumea family members with $H_{r}>5$ are not corrected, as the OSSOS survey reveals a significant scarcity of small members within the Haumea family (Pike et al., 2020)., based on the OSSOS++ survey (Petit et al., 2023). Their Figure 1 presents the differential distribution, which we fit by a broken power law
@@ 2.2 Growth of bodies by pebble accretion
where $\chi_{\bullet}$ is the physical-to-Hill radius ratio
@@
where $\rho_{\bullet}$ the internal density of the planetary body and $A_{\mathrm{geo}}=\pi/3^{2/3}\approx 1.5$. In addition, the characteristic growth timescale is
@@
where $\rho_{\mathrm{peb}}=Z_{\mathrm{peb}}\rho_{\mathrm{gas}}=Z_{\mathrm{peb}}\Sigma_{\mathrm{gas}}/\sqrt{2\pi}H_{\mathrm{gas}}$ the mass density of pebbles in the midplane, $Z_{\mathrm{peb}}$ the corresponding local metallicity, $\Sigma_{\mathrm{gas}}$ the surface density of the gas and $H_{\mathrm{gas}}$ the gas pressure scale height. The disk aspect ratio $h=H_{\mathrm{gas}}/r$ can also be written
@@
The degree of coupling between particles and gas is quantified by the particle stopping time $t_{\mathrm{stop}}$, which represents the time for a particle to adjust its motion to that of gas. Its dimensionless equivalent, $\tau_{s}=\Omega_{K}t_{\mathrm{stop}}$, is referred to as the aerodynamic size (in the literature, the Stokes number or St, is also common). In this work we define pebbles as particles of $\tau_{s}<1$. For the Epstein drag, $\tau_{s}$ can be expressed as $\tau_{s}=\pi(\rho_{\bullet}R)_{\mathrm{peb}}/2\Sigma_{\mathrm{gas}}$, indicating that the aerodynamic size is proportional to $R_{\mathrm{peb}}$.
The relative velocity $\Delta v=\sigma v_{K}$ can be identified in one of three ways:
  • the eccentricity of planetesimals, $\Delta v\simeq ev_{K}$ ($\sigma=e$), which is not considered in this work;
@@
    (e.g., Weidenschilling, 1977) is a parameterization of the radial pressure gradient. Pebbles with $\tau_{s}<1$ follow the gas, while planetesimals ($\tau_{s}\gg 1$) move at the Keplerian velocity. Hence, the pebble-planetesimal velocity is also $\eta v_{K}$;
@@
Roughly, the highest of these velocities ($e$, $\eta$, or $\mathcal{M}h$) determines $\Delta v$.
Gravitational focusing increases the accretion rate by a factor of $(v_{\mathrm{esc}}/\Delta v)^{2}=2Gm/R(\Delta v)^{2}$, where $v_{\mathrm{esc}}$ is the escape velocity from its surface. In dimensionless units, the rate becomes
@@
Gravitational focusing significantly enhances the growth rates of planetesimals. On the other hand, its efficacy is hampered by the more copious noncollisional scatterings of planetesimals – a process known as viscous stirring (Ida & Makino, 1993) – or turbulent stirring (Ida et al., 2008), which increases $\Delta v$. For small pebbles tightly coupled to the gas, this is not a concern, because gas drag circularizes their orbits before the next encounter. Therefore, in contrast to planetesimals, there is no need to follow the evolution of $\Delta v$ with time.
More potently, pebbles accrete by the settling mechanism, which is better known as pebble accretion (Ormel & Klahr, 2010; Lambrechts & Johansen, 2012). In pebble accretion, pebbles can be captured at distances much larger than $R$, after which the pebbles settle down in the gravitational potential. We adopt the numerically calibrated expression by Ormel & Liu (2018) [^6] Ormel & Liu (2018) express these expression in terms of an efficiency $\epsilon=\dot{M}/2\pi r\Sigma v_{r}$. We take their 3D expression, $\epsilon=0.39q/\eta h_{\mathrm{peb}}\times f_{\mathrm{set}}^{2}$. With the drift velocity $v_{r}=2\eta v_{K}\tau_{s}$ and $\rho_{\mathrm{peb}}=\Sigma_{\mathrm{peb}}/\sqrt{2\pi}h_{\mathrm{peb}}$, we obtain Equation (10).:
$\frac{dq}{dt}=A_{\mathrm{peb}}\frac{q\tau_{s}}{t_{0}}f_{\mathrm{set}}^{2}.$ (10)
Here $A_{\mathrm{peb}}=12.3$ and $f_{\mathrm{set}}\leq 1$ is a modulation factor that suppresses accretion rates at low masses. It is a defining characteristic of pebble accretion that rates do not depend on the physical radius (or the nondimensional $\chi_{\bullet}$ parameter; Ormel 2017). In addition, we assumed the 3D accretion regime for pebbles in Equation (10), for which $\sigma$ does not enter explicitly. The velocity matters in the modulation factor $f_{\mathrm{set}}$, however:
$f_{\mathrm{set}}^{2}=\exp\left[-\left(\frac{\Delta v}{v_{\ast}}\right)^{2}\right]\quad\textrm{or}\quad f_{\mathrm{set}}^{2}=\left[1+a_{\mathrm{turb}}\left(\frac{\Delta v}{v_{\ast}}\right)^{2}\right]^{-3}.$ (11)
Here $v_{\ast}=(q_{p}/\tau_{s})^{1/3}v_{K}$ and $a_{\mathrm{turb}}=0.33$ (Liu & Ormel, 2018). The first expression in Equation (11) holds for a laminar-dominated velocity field and the second for a turbulent-dominated $\Delta v$. In the case of a turbulent flow, $\Delta v$ should be regarded as the rms value of an underlying velocity distribution that is isotropic (Ormel & Liu, 2018).